OpenStudy (anonymous):

the area of a circle is 100 pi sq. in. What is the circumference

hero (hero):

Hint: \[A_{circle} = \pi r^2\]\[C_{circle} = 2 \pi r\] Input known value for Area, solve for r, substitute r into circumference formula